OpenStudy (anonymous):

how many solutions does 3x-7=2x-1+x+7 have?

OpenStudy (anonymous):

3x-7=2x-1+x+7 3x - 2x - x = -1 + 7 + 7 0 = - 1 + 7 + 7 because of 3x - 2x - x = 0, then 3x - 2x - x = 0 (3-2-1)x = 0 x = 1 , or x = 0

jhonyy9 (jhonyy9):

so i think will be 3x-7=2x-1+x+7 what will be 3x-2x-x = 7-1+7 0 = 13 so do you know what mean this when will be eliminated the unknowed term x for what you need to solve this equation ? so this mean that this equation have not solution(s) because indifferent what value(s) you will add to x never not will get the laft side equal right side hope this is understandably sure right easy good luck bye